# Agentic Program Verification

### v1

@@ 8. Related Work
8.1. Formal Verification Tools and Practices
Our work continues the advances on forging an end-to-end verification workflow for software, and thus is closely related to pioneering efforts on verifying critical software systems (DBLP:conf/sosp/KleinEHACDEEKNSTW09; DBLP:conf/osdi/ChajedTTKZ22; leroy2016compcert; verified_db), network protocols (DBLP:conf/sosp/ZhengLXSLHWMLCZ23), and microprocessor designs (DBLP:journals/dt/JonesOSAM01). Automation in formal verification is largely underpinned by verification-oriented languages including Frama-C (DBLP:journals/fac/KirchnerKPSY15), Dafny (DBLP:conf/lpar/Leino10), Verus (DBLP:journals/corr/abs-2303-05491), and Viper (DBLP:conf/vmcai/0001SS16), which integrate specification and proof directly into the programming workflow through the use of annotations. These tool chains generate proof obligations automatically and then attempt to discharge them to automated solvers (DBLP:conf/tacas/MouraB08; barbosa2022cvc5). When such automation is insufficient, interactive theorem provers (ITPs), such as Isabelle/HOL (paulson1994isabelle), Lean (DBLP:conf/cade/Moura021), and Rocq (coq1996coq), are used as the back end of the workflow. Our approach critically depends on the verification infrastructure.
8.2. Machine Learning for Program Verification
Our work builds on a rapidly growing body of work (lisurvey) that applies machine learning to program verification. Related approaches include loop invariants inference (DBLP:conf/cav/SiNDNS20; si2018learning), automatic insertion of annotations (DBLP:journals/tmlr/LoughridgeSACS025), and direct synthesis of verified methods (DBLP:journals/pacmse/MisuLM024). Among these directions, our work is most closely related to automatic approaches in theorem proving. We introduce a new paradigm in this space: agentic approaches, in which agents interpret partial proofs and collaborate with a theorem prover to construct complete ones. Our method extends and unifies three broad paradigms of prior work, as detailed below.
Neural Proof Generation. A significant body of work treats theorem proving as a sequence-to-sequence (seq2seq) translation task, and thus employs neural networks to learn from large corpora of human-written proofs (yang2019coqgym; sanchez2020generating; first2020tactok). While effective at capturing common proof patterns, these networks generate tactics purely based on their parametric knowledge, without a fine-grained, dynamic understanding of the evolving proof state. This limitation is particularly acute in software verification, where proofs are long, the state space is enormous, and the correct next step is highly sensitive to the precise context.
Retrieval-Augmented Proving. To augment parametric models with more context, recent works have integrated retrieval mechanisms to fetch relevant premises from a large library of existing theorems during proof generation (lu2025palm; rango-icse25). This approach is promising because it grounds the generation process in established facts. However, these approaches perform retrievals statically, lacking a tight, iterative feedback loop between the proof state, the retriever, and the generator. Consequently, the retriever may fetch axioms that are relevant to the overall theorem but useless for the specific sub-goal the prover is currently tackling.
Reinforcement Learning. To better understand the proving progress, approaches like (DBLP:conf/nips/WuNWD21) employ reinforcement learning, where each tactic application is awarded with immediate feedback to facilitate incremental progress. The inherent sparsity of positive feedback prompts clever design of reward functions that estimate the value of intermediate states (sanchez2024qedcartographer). However, such reward functions often provide noisy signals or overfit to short-term goals, thereby hindering strategic, long-horizon reasoning.
@@ 9. Perspectives
As AI-generated code becomes increasingly prevalent in software development, the need for automated verification becomes more pressing. Our work demonstrates that LLM agents can bridge the gap between automatic code generation and formal verification, moving closer to the vision of trusted automatic programming. The agentic nature of our approach suggests a paradigm shift where verification tools act as intelligent collaborators rather than passive validators, capable of autonomously navigating complex proof spaces and adapting to diverse verification challenges. This represents a crucial step toward making formal verification accessible for real-world software systems, where the combination of automated reasoning and agentic intelligence can provide the rigor and scalability needed for next-generation software assurance.

### v3

@@ 8. Related Work
8.1. Deductive Program Verification
Deductive verification formally verifies the correctness of software systems by extracting and proving a collection of mathematical proof obligations from both the program and its specification (Filliâtre and Marché, 2007). The automation of the verification process is largely underpinned by verification-oriented languages, including Frama-C (Kirchner et al., 2015), Dafny (Leino, 2010), Verus (Lattuada et al., 2023), and Viper (Müller et al., 2016), which provide an end-to-end verification flow that integrates specification and proof directly into the programming workflow through the use of annotations. These tool chains generate proof obligations automatically, thereby reducing the problem of program verification to theorem proving. The proof obligations are sent to either Automated Theorem Provers (ATPs) or Interactive Theorem Provers (ITPs) to verify. ATPs such as CoqHammer (Czajka and Kaliszyk, 2018) attempt to discharge the obligations automatically through constraint solving (De Moura and Bjørner, 2008; Barbosa et al., 2022), while ITPs such as Isabelle/HOL (Paulson, 1994), Lean (de Moura and Ullrich, 2021), and Rocq (The Coq Development Team, 2023) allow users to interactively construct proofs using tactics. Due to ATPs’ limited capabilities in handling complex proofs (Nieuwenhuis et al., 2007) and occasional correctness issues (Mansur et al., 2020; Winterer et al., 2020), ITPs have become the de facto standard for verifying sophisticated software systems. Indeed, many pioneering efforts have successfully verified critical software systems (Klein et al., 2009; Chajed et al., 2022; Leroy et al., 2016; Malecha et al., 2010), network protocols (Zheng et al., 2023), and microprocessor designs (Jones et al., 2001) with ITPs. However, ITPs require significant human effort to guide the proof construction, which limits their scalability and applicability. AutoRocq specifically focuses on reducing the manual effort in proving complex residual obligations, thus advancing towards an automatic, end-to-end verification workflow. As such, it is related but complementary to prior work automating other aspects of the workflow, including inference of specifications (Endres et al., 2024) and loop invariants (Si et al., 2020, 2018; Loughridge et al., 2025; Garg et al., 2016), discovery of helper lemmas (Sivaraman et al., 2022), and direct synthesis of verified methods (Misu et al., 2024).
8.2. Automated Techniques in Theorem Proving
Our work is closely related to the substantial research devoted to automating mathematical theorem proving (Li et al., 2024b; Google DeepMind, 2024), spanning the complete pipeline from premise selection and tactic generation to proof search and recovery mechanisms. Our focus on proving program-derived lemmas leads to unique design choices in AutoRocq, which we detail below.
Premise Selection. Premise selection is the task of identifying relevant lemmas, definitions, or axioms from a large library to assist in proving a given theorem (Yang et al., 2023; Mikuła et al., 2023). Early approaches treat premise selection as a binary classification task, where the goal is to predict whether a given premise is relevant to the theorem at hand (Alemi et al., 2016). More recent approaches such as CoqHammer (Czajka and Kaliszyk, 2018), PALM (Lu et al., 2024), and Rango (Thompson et al., 2025) leverage machine learned weights or lexical similarity measures (Sparck Jones, 1972) to rank premises statically. In contrast, AutoRocq adopts on-demand premise retrieval with the support of pattern-based context search.
Tactic Generation. Tactic generation is the core component in interactive theorem proving, where a proof step is proposed to transform the current proof state (Li et al., 2024b). Early works rely on hand-crafted heuristics and domain-specific knowledge to guide generation (Leroy et al., 2016; Gu et al., 2016; Klein et al., 2009). More recently, machine learning techniques have been applied to learn step-wise generation strategies from large corpora of human-written proofs (Wu et al., 2021; Google DeepMind, 2024; Blaauwbroek et al., 2020; Yang and Deng, 2019; Sanchez-Stern et al., 2020; First et al., 2020; Thompson et al., 2025). AutoRocq follows the same paradigm and queries the underlying LLM at each step with a tree-based proof state representation.
Proof Search. Proof search concerns the sequence of tactic applications, and has been a long-standing challenge due to an inherently large search space and sparse rewards (Czajka, 2020; Lample et al., 2022). Whole-proof generation approaches such as PALM (Lu et al., 2024) circumvent the search process by predicting the entire tactic sequence in a single pass. Step-by-step proving techniques conventionally employ beam search to sample multiple tactic predictions with breadth-first (Bansal et al., 2019), depth-first (Yang and Deng, 2019), or best-first heuristics (Polu and Sutskever, 2020; Thompson et al., 2025) to traverse the search space. Recent works have leveraged other exploration mechanisms such as reinforcement learning (Wu et al., 2021; Sanchez-Stern et al., 2025) and in-context learning (Thakur et al., 2024). AutoRocq adapts the search strategy through its agency, enabled by timely feedback, high-level interpretation of the proving progress, and on-the-fly learning from successful histories.
Recovery Mechanisms. Due to the trial-and-error nature of theorem proving, proof automation systems need to effectively recover proof attempts and explore alternative strategies (Lu et al., 2024; Thakur et al., 2024). Yet this issue is often not addressed explicitly. Rather, a large body of work takes only binary signals from ITPs and simply retries tactic generation in case of failures (Thompson et al., 2025; Sanchez-Stern et al., 2025). Others, such as Copra (Thakur et al., 2024) and Proverbot9001 (Sanchez-Stern et al., 2020), are able to restore to an earlier proof state if necessary. PALM (Lu et al., 2024), on the other hand, performs opportunistic tactic repairs and delegates the open subgoal to an ATP (Czajka and Kaliszyk, 2018) as the last resort. In contrast, AutoRocq exploits detailed diagnostic errors from ITPs to rectify failed tactic applications, and handles persistent errors through autonomous context enrichment.
8.3. LLM for Software Quality Assurance
Besides formally verifying correctness through theorem proving as is done in AutoRocq, a wide spectrum of techniques for software quality assurance have benefited from recent advances in Large Language Models (Hou et al., 2024; Fan et al., 2023). For example, model checking techniques also offer strong correctness guarantees by encoding a software system as a finite state machine and iteratively checking its states and transitions. LLMs have been employed to automate the construction of both the model (Zuo et al., 2025) and the formal specification (Wu et al., 2024a). Testing techniques, on the other hand, try to prove the presence of bugs by synthesizing test cases that trigger unintended behaviors, and have benefited significantly from LLMs’ capability in input generation (Meng et al., 2024; Deng et al., 2024; Xia et al., 2024) and program understanding (Tu et al., 2026; Luo et al., 2026). Static analysis leverages LLMs to identify code issues and vulnerabilities faster and earlier in the development process (Li et al., 2024a; Wu et al., 2024b). Recent work also applies agentic hybrid (static–dynamic) analysis to security tasks such as cross-platform rule conversion (Xu et al., 2026). While each technique above presents unique trade-offs, deductive verification stands out for its mathematical rigor and stringent guarantees on correctness, making it particularly well-suited for safety-critical systems (Efremov et al., 2018).
@@ 9. Perspectives
As AI-generated code becomes increasingly prevalent in software development, the need for automated verification becomes more pressing. Our work demonstrates that LLM agents can bridge the gap between automatic code generation and formal verification, moving closer to the vision of trusted automatic programming. The agentic nature of our approach suggests a paradigm shift where verification tools act as intelligent collaborators rather than passive validators, capable of autonomously navigating complex proof spaces and adapting to diverse verification challenges. This represents a crucial step toward making formal verification accessible for real-world software systems. We can thus move closer to the vision of AI-based Verification and Validation (V & V) of AI-generated code. This will alleviate the problem of humans reviewing AI-generated code.
Acknowledgements.
We would like to sincerely thank all the anonymous reviewers for their valuable comments. This work was partially supported by a research project “AI for Program Reasoning”, project AI4SCH-2025-0084 under AI for Science program of National Research Foundation (NRF) Singapore. Any opinions, findings and conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.
Data Availability
We release AutoRocq, including its implementation, benchmarks, and replication instructions, for academic use at the following link: https://github.com/NUS-Program-Verification/AutoRocq.